OpenStudy (anonymous):

A rock is thrown upward with a velocity of 28 meters per second from the top of a 23 meter high cliff, and it misses the cliff on the way back down. When will the rock be 4 meters from the water, below?

OpenStudy (anonymous):

Will some one please help me

OpenStudy (whpalmer4):

Do you know the general formula\[h(t) = -\frac{1}{2}gt^2 + v_0t + h_0\]for the height of an object launched with initial velocity \(v_0\) and initial height \(h_0\)?

OpenStudy (whpalmer4):

Set that equal to the height of the rock when it is 4 meters from the water and solve for the positive value of \(t\).

OpenStudy (whpalmer4):

\[g=9.8\text{ m/s}^2\]

OpenStudy (anonymous):

i am sorry the numbers are 11 meters per second and 24 meter high cliff and when will the rock be 8 meters from the water

OpenStudy (anonymous):

i have no idea how to do this

OpenStudy (whpalmer4):

Exactly the same way as if the numbers were 28 m/s, 23 meter high cliff, 4 meters from the water.

OpenStudy (anonymous):

ok i don't know how to set it up or solve it, i am really bad at algebra

OpenStudy (whpalmer4):

Well, baby steps. First, identify the variables that correspond to your numbers.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

h=8 g=9.8 v0=11m/s h0=24

OpenStudy (anonymous):

right?

OpenStudy (whpalmer4):

yes!

OpenStudy (anonymous):

ok then what?

OpenStudy (whpalmer4):

so write out the formula

OpenStudy (anonymous):

well there are t's in the formula but i don't have number for those

OpenStudy (whpalmer4):

Yes, we need to have a variable to solve for :-)

OpenStudy (whpalmer4):

so just write \(t\) in the appropriate spots for \(t\)

OpenStudy (anonymous):

8=1/2(9.8)t^2+11t+8

OpenStudy (anonymous):

right?

OpenStudy (whpalmer4):

Nope. Go back and look at my equation carefully. There are at least two errors in your formula

OpenStudy (anonymous):

h=-1/2

OpenStudy (anonymous):

but im not sure where the other error is?

OpenStudy (whpalmer4):

Also, you should put the 1/2 as (1/2) to make it crystal clear that it is \[-\frac{1}{2}gt^2\]and not\[-\frac{1}{2gt^2}\]

OpenStudy (whpalmer4):

Okay, the missing negative sign was one error. Can you please type the whole equation as you think it should be now?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

8=(-1/2)(9.8)t^2+11t+8

OpenStudy (anonymous):

right?

OpenStudy (whpalmer4):

No. I'm curious about the two 8's in the formula. Wouldn't they just cancel each other out?

OpenStudy (whpalmer4):

Can you tell me why 8 appears twice in the formula?

OpenStudy (anonymous):

well i was kind of confused with the h0 and just the h

OpenStudy (whpalmer4):

okay. let's look at my original formula: \[h(t) = -\frac{1}{2}gt^2+v_0t+h_0\]\(h(t)\) is the height at time \(t\) \(h_0\) is the initial height, as I mentioned when I wrote the original formula Which, if either, of those should be 8 in your formula to find the time at which the rock is 8 meters above the surface of the water?

OpenStudy (anonymous):

i don't think it should be in the equation? right

OpenStudy (whpalmer4):

What is the initial height of the rock?

OpenStudy (anonymous):

0?

OpenStudy (whpalmer4):

Isn't it being thrown off a cliff?

OpenStudy (anonymous):

yah

OpenStudy (anonymous):

24meters high

OpenStudy (whpalmer4):

Okay, so \[h_0 = 24\]What do we have for a formula now?

OpenStudy (anonymous):

8=(-1/2)(9.8)t^2+11t+24

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

now what?

OpenStudy (whpalmer4):

Yes. Now use the quadratic formula to solve that for \(t\). If \[ax^2+bx+c=0,\,a\ne0\]\[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

−0.561224489796+1.125459783873i,−0.561224489796−1.125459783873i

OpenStudy (anonymous):

so how many secs is that?

OpenStudy (whpalmer4):

no, you must have done the arithmetic incorrectly. what did you use for a, b, c?

OpenStudy (anonymous):

i don't know my teacher gave me a website to type in a an equation and it tells you the answer for the quadric formula

OpenStudy (anonymous):

and it is usually right

OpenStudy (whpalmer4):

what did you type in?

OpenStudy (whpalmer4):

I'm betting it is always right if you type the formula correctly :-)

OpenStudy (anonymous):

8=(-1/2)(9.8)t^2+11t+24

OpenStudy (anonymous):

bc i have put stuff in there like that and it gives me the right answer

OpenStudy (anonymous):

what is it suppose to be

OpenStudy (whpalmer4):

well, for starters, that isn't in the appropriate form, is it? I said the quadratic formula applies if you have a function \[ax^2+bx+c=0\] Is your equation in that form? what is the url of this website?

OpenStudy (whpalmer4):

I must say I'm not in favor of students not knowing how to solve quadratic equations with the formula without computer assistance!

OpenStudy (anonymous):

well how do i get in to that format?

OpenStudy (whpalmer4):

Using the computer to do the drudge work once you've learned how to do it yourself is a completely different matter. Do you have everything on one side, and 0 on the other? If not, do some basic algebra until you do.

OpenStudy (anonymous):

0=(-1/2)(9.8)t^2+11t+24-8

OpenStudy (whpalmer4):

yes. Can you simplify that at all?

OpenStudy (anonymous):

0=(-4.9)t^2+11t+16

OpenStudy (whpalmer4):

Good. so what are the values of \(a,b,c\) to go into the quadratic formula?

OpenStudy (anonymous):

a=-4.9t^2 b=11t c=16

OpenStudy (whpalmer4):

no, \(a,b,c\) are just the coefficients, the variable doesn't come along for the right

OpenStudy (anonymous):

what????

OpenStudy (whpalmer4):

if \(a=-4.9, b=11, c=16\), then what is \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]just write it out first and show me, don't try to evaluate it

OpenStudy (whpalmer4):

"come along for the ride" was what I meant to type